OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region under the curve y = 3 - x about the x-axis over the interval [0, 1].

OpenStudy (amistre64):

you do realize that this makes a cone right?

OpenStudy (amistre64):

well, cone like :) it appears to be cut off at 1 other than 3

OpenStudy (anonymous):

volume= 2pi (int(3-x) from 0 to 1)

OpenStudy (amistre64):

they give a polynomial to build your confidence with since they are the simplest things to integrate

OpenStudy (amistre64):

an no, the volume of rotation is adding up areas, not circumferences ... so it aint gonne a be a 2pi[f(x)]

OpenStudy (amistre64):

to add up all the areas of the circles created from 0 to 1; we have to integrate pi [f(x)]^2

OpenStudy (amistre64):

\[\pi\ * \int_{0}^{1}\ (3-x)^2\ dx\] \[\pi\ * \int_{0}^{1}\ (9-6x+x^2)\ dx\] \[\pi\ \left( \int_{0}^{1}\ (9)dx -\int_{0}^{1}(6x)dx+\int_{0}^{1}(x^2)\right)\] \[\pi\ \left( 9\int_{0}^{1}dx -6\int_{0}^{1}(x)dx+\int_{0}^{1}(x^2)\right)\] \[\pi\ \left( 9x|^{1} -6\frac{x^2}{2}|^{1}+\frac{x^3}{3}|^{1}\right)\] \[\pi\ \left( 9 -3+\frac{1}{3}\right)\]

OpenStudy (amistre64):

missed a few dxs in there but it should still be readable :)

OpenStudy (anonymous):

So the final answer is (19/3)π right?

OpenStudy (amistre64):

dunno, I forgot how to add and multiply :) too much calculus to remember ....

OpenStudy (amistre64):

19/3 pi sounds good